A buffer solution is prepared by mixing 78.0 mL of 0.279 M fluorobenzoic acid (C7H5O2F) with 50.0 mL of 0.476 M sodium fluorobenzoate. The pKa of this weak acid is 2.90. What is the pH of the buffer solution? (with correct sig figs)
WebAssign will check your answer for the correct number of significant figures.  
Answer #1

Answer

pH = 2.94

Explanation

After mixing the solution

Concentration of C7H5O2F = 0.279M/( 128ml/78ml) = 0.1700M

Concentration of C7H4O2F- = 0.476M/( 128ml/ 50ml) = 0.1859M

Applying Henderson - Hasselbalch equation

pH = pKa + log([A-]/[HA])

pH = pKa + log( [ C7H4O2F-]/[C7H5O2F])

pH = 2.90 + log( 0.1859M/0.1700M)

pH = 2.90 + 0.039

pH = 2.94